# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ Abstract
We propose an iterative algorithm for clustering high-dimensional data, where the true signal lies in a much lower-dimensional space. Our method alternates between feature selection and clustering, without requiring precise estimation of sparse model parameters. Feature selection is performed by thresholding a rough estimate of the discriminative direction, while clustering is carried out via a semidefinite programming (SDP) relaxation of K-means. In the isotropic case, the algorithm is motivated by the minimax separation bound for exact recovery of cluster labels using varying sparse subsets of features. This bound highlights the critical role of variable selection in achieving exact recovery. We further extend the algorithm to settings with unknown sparse precision matrices, avoiding full model parameter estimation by computing only the minimally required quantities. Across a range of simulation settings, we find that the proposed iterative approach outperforms several state-of-the-art methods, especially in higher dimensions.
@@ 1 Introduction
High-dimensional clustering algorithms group unlabeled observations when the number of features exceeds or matches the sample size. It is widely applied to data from advanced recording technologies, such as neurophysiological recordings (Cunningham and Yu,, 2014; Kadir et al.,, 2014; Park et al.,, 2024), gene expression (Bittner et al.,, 2000; Curtis et al.,, 2012; He et al.,, 2021), and neuroimaging (Eshaghi et al.,, 2021; Heller et al.,, 2006; Poulakis et al.,, 2022), which produce high-dimensional measurements but limited samples due to acquisition costs. In these settings, typically only a small subset of features is relevant for clustering. Then, a key challenge lies in the interdependence between the two unsupervised tasks of variable selection and clustering. On one hand, with unknown sparsity, including too many noise features in the clustering algorithm hinders the results. On the other hand, it is challenging to pinpoint the signal features when the true cluster structure is unknown. This motivates an iterative algorithm that alternates between clustering and variable selection until convergence leading to the identification of the true clusters.
Sparse high-dimensional clustering model.
To effectively study the role of sparsity, we assume a Gaussian mixture model with sparse discriminative directions. Let $\mathbf{X}\in\mathbb{R}^{p\times n}$ denote the data matrix, where $p$ may exceed $n$. Its columns $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}$ are independently generated as $\mathbf{X}_{i}\sim\mathcal{N}(\boldsymbol{\mu}^{\ast}_{k},\boldsymbol{\Sigma}^{\ast})$ if $i\in G_{k}^{\ast}$, where the non-random sets $G_{1}^{\ast},\ldots,G_{K}^{\ast}$ form a partition of $\{1,\ldots,n\}$. The number of clusters $K$ is fixed and known. The unknown cluster centers are denoted by $\boldsymbol{\mu}^{\ast}_{1},\ldots,\boldsymbol{\mu}^{\ast}_{K}\in\mathbb{R}^{p}$, and $\boldsymbol{\Sigma}^{\ast}\in\mathbb{R}^{p\times p}$ is the common covariance matrix. Let us denote the true precision matrix as $\boldsymbol{\Omega}^{\ast}:=(\boldsymbol{\Sigma}^{\ast})^{-1}$. We consider both cases where $\boldsymbol{\Omega}^{\ast}$ is known and where it is unknown but have a sparse structure. Our objective is to recover $G_{1}^{\ast},\dots,G_{K}^{\ast}$ while treating $\boldsymbol{\mu}^{\ast}_{k}$’s and $\boldsymbol{\Sigma}^{\ast}$ as nuisance parameters. We assume that only a small, unknown subset of features, indexed by $S_{0}$, is relevant for clustering. Formally, $S_{0}:=\bigcup_{1\leq k\neq l\leq K}\mathrm{supp}\bigl{(}\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})\bigr{)}\subset\{1,\ldots,p\}$ with $|S_{0}|\ll p$. This assumption aligns with results from classification of two-class Gaussian mixtures, where the optimal decision boundary is proportional to $\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$, maximizing the signal-to-noise ratio (Fan et al.,, 2013).
Previous methods.
@@
High-dimensional sparse clustering methods roughly fall into two types: two-step and iterative. We start with the two-step type. Since clustering after conventional dimension reduction, such as PCA (Ghosh and Chinnaiyan,, 2002; Liu et al.,, 2003; Tamayo et al.,, 2007), can miss a sparse clustering signal (Chang,, 1983), many methods adopt (Gaussian mixture) models to select features based on the clustering signal. They select features using multiple testing based on goodness-of-fit (Jin and Wang,, 2016) or moments (Azizyan et al.,, 2013; Verzelen and Arias-Castro,, 2017), model selection (Maugis et al.,, 2009; Raftery and Dean,, 2006), sparse PCA (Löffler et al.,, 2022), or an estimated sparse discriminating direction (denoted $\boldsymbol{\beta}^{\ast}$; see Section 3.1) (Azizyan et al.,, 2015). More examples include Bouveyron and Brunet-Saumard, (2014); Brodinová et al., (2019); Kadir et al., (2014); Liu et al., (2023). Since it is difficult to identify the signal features without knowing the cluster structure, two-step methods often require a stronger signal compared to iterative methods (compare the conditions in Jin and Wang, (2016) with Cai et al., (2019)). On the other hand, iterative methods repeatedly refine cluster assignments and model parameters estimates. Most methods alternate between clustering and regularized estimation of model parameters, such as feature weights (Chakraborty et al.,, 2023; Huo and Tseng,, 2017; Li et al.,, 2018; Tsai and Chiu,, 2008; Witten and Tibshirani,, 2010; Yuan et al.,, 2024), cluster centers (Sun et al.,, 2012), and feature rankings (Arias-Castro and Pu,, 2017; Zhang et al.,, 2020). Regularized EM algorithms follow the same principle: the E-step updates cluster assignments, while the M-step performs regularized estimation of sparse model parameters (Bouveyron et al.,, 2007; Dong et al.,, 2024; Fraley and Raftery,, 2007; Fu et al.,, 2021; Guo et al.,, 2010; Pan and Shen,, 2007; Wang and Zhu,, 2008; Xie et al.,, 2008; Zhou et al.,, 2009). Notably, Cai et al., (2019) establish a minimax rate of order $(s\log p)/n$ for the excess clustering risk on new samples, based on estimating $\boldsymbol{\beta}^{\ast}$ from the training data using $\ell_{1}$-regularized EM. Sparse Bayesian hierarchical models (Tadesse et al.,, 2005; Yao et al.,, 2025) and sparse convex clustering (Chakraborty and and Xu,, 2023; Wang et al.,, 2018) also follow the iterative paradigm, as their fitting procedures, such as MCMC or ADMM , progressively refine parameter estimates and feature selection.
Our contributions.
In the existing literature of iterative methods, clustering heavily depends on accuracy of estimated nuisance parameters such as cluster centers. Additionally, their estimation requires tuning sparsity-inducing regularization parameters or hyperpriors, which are challenging to tune in unsupervised problems. In contrast, our method alternates between clustering and feature selection, where our feature selection step does not require precise estimation of nuisance parameters. Features are selected by thresholding a rough estimate of the discriminating direction $\boldsymbol{\beta}^{\ast}$; this estimate is used solely to guide feature selection and has no direct impact on the clustering step. Based on the selected features, clustering is carried out using an SDP-relaxed K-means on a transformed dataset, avoiding accurate estimation of cluster centers or the covariance matrix. To obtain the transformed dataset, we adapt a high-dimensional precision matrix estimation tool (Fan and Lv,, 2016), modifying it to estimate only the minimally required quantities without estimating high dimensonal nuisance parameters. In the case of a known isotropic covariance matrix, we motivate our algorithm by deriving a minimax separation bound for exact recovery under varying feature subsets by SDP $K$-means, which scales as $\log n+s\log p/n+\sqrt{s\log p/n}$, while the minimax lower bound scales as $\log n$. In the common high-dimensional regime where $s\log p/n\leq\log n$, the upper and lower bounds match.
Notation.
@@
We denote vectors and matrices by boldface letters, and their individual elements by the corresponding non-boldface letters. For a matrix $\mathbf{A}$, single and double subscripts denote columns and (row, column) elements, respectively. A subscript may be a single index or an index set. For vectors not derived from a matrix, a subscript denotes an element or a sub-vector. If the vector or matrix already contains a subscript, we use parentheses to avoid ambiguity, for example, $(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{S}$. We write $\mathbf{A}\geq 0$, $\mathbf{A}>0$, and $\mathbf{A}\succeq 0$ to denote elementwise nonnegativity, elementwise positivity, and positive semidefiniteness of a matrix, respectively. For a set $S$, let $|S|$ denotes its cardinality. For a positive integer $a$, let $\mathbf{1}_{a}\in\mathbb{R}^{a}$ denote the all-one vector, and $[a]$ denote $\{1,\cdots,a\}$. For a set $G_{k}\subset[n]$, let $\mathbf{1}_{G_{k}}\in\{0,1\}^{n}$ denote the indicator vector that takes the value 1 at indices in $G_{k}$ and 0 elsewhere.
2 SDP relaxation of K-means and theoretical motivation
Section 2.1 presents the SDP relaxation of $K$-means (Peng and Wei,, 2007; Zhuang et al.,, 2023) and its limitation in high dimensions due to its inability to adapt to sparsity, motivating our sparsity-aware extension. Section 2.2 provides theoretical motivation for the proposed approach, supporting iterative methods that solve multiple SDP relaxations over varying subsets of variables and highlighting the importance of sparsity.
2.1 SDP relexation of K-means problem and its limitation in high dimensions
@@
It has been derived in (Peng and Wei,, 2007; Zhuang et al.,, 2023) that finding the partition $G_{1},\ldots,G_{K}$ that maximizes the K-means objective can be formulated as the following NP-hard mixed-integer program (Mahajan et al.,, 2012):
$\max_{\mathbf{H}\in\{0,1\}^{n\times K}}\;\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{H}\mathbf{B}\mathbf{H}^{\top}\rangle,\quad\text{s.t.}\quad\mathbf{H}\mathbf{1}_{K}=\mathbf{1}_{n},$ (1)
@@
where the inner product denotes the Frobenius inner product, $\mathbf{B}:=\mathrm{diag}(|G_{1}|,\ldots,|G_{K}|)$, and each row of $\mathbf{H}$ is a one-hot cluster indicator, formally $H_{i,k}:=\mathbbm{1}(\mathbf{X}_{i}\in G_{k})$ for $i\in[n]$ and $k\in[K]$. This objective is connected to the log likelihood function in the Gaussian mixture model (see Appendix E.1 for details). By observing $\mathbf{Z}=\mathbf{H}\mathbf{B}\mathbf{H}^{\top}$ satisfies $\mathbf{Z}^{\top}=\mathbf{Z},\mathbf{Z}\succeq 0,\mathrm{tr}(\mathbf{Z})=K,\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n}$, and $\mathbf{Z}\geq 0$, we lift $\mathbf{Z}$ into the convex space and solve the following semidefinite problem:
@@
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{Z}\rangle,\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~{}\mathbf{Z}\succeq 0,~{}\mathrm{tr}(\mathbf{Z})=K,~{}\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~{}\text{and}~{}\mathbf{Z}\geq 0.$ (2)
This relaxation, derived in (Zhuang et al.,, 2023), can be solved using off-the-shelf SDP solvers (for example, (Sun et al.,, 2020)). Moreover, faster approximation methods for solving this relaxation have been proposed (Zhuang et al.,, 2022, 2024).
@@
When $\boldsymbol{\Sigma}^{\ast}$ is known to be $\sigma^{2}\mathbf{I}_{p}$ with fixed dimensionality $p$, Chen and Yang, (2021) proves that this SDP relaxation is a minimax optimal clustering algorithm under the exact recovery loss, with the minimax separation rate for clustering scaling as $\sigma^{2}\big{(}\log n+\sqrt{(\log n)^{2}+Kp/n}\big{)}$. This rate increases with $p$ at a square root rate. It also highlights that, the SDP relaxation may fail in high-dimensional settings where $p$ is comparable or even larger than $n$, unless the separation between clusters also scales with $p$ quickly at a square root rate – an assumption that can be very stringent for large $p$.
On the other hand, sparsity assumption has been commonly imposed in the literature to address the challenges related to high dimensionality, and has been verified to hold true in broad application scenarios such as medical imaging (Mishra et al.,, 2017), medical signal processing (Zhang et al.,, 2021), financial modeling (Nystrup et al.,, 2021), and genetics (Lee et al.,, 2015). This motivates us to investigate how to modify the SDP $K$-means algorithm to incorporate sparsity in high dimensions and explore the corresponding separation rate. Before presenting our algorithm, we first theoretically investigate the fundamental difficulty of the problem and understand the role of sparsity in high-dimensional clustering using SDP relaxation.
2.2 Minimax rate for separation bound in isotropic case
For theoretical study, this section assumes that $\boldsymbol{\Sigma}^{\ast}=\sigma^{2}\mathbf{I}_{p}$ is known. We study how the separation bound for exact recovery scales with dimension $p$, sample size $n$, and sparsity $s$. We adopt a minimax framework to assess worst-case performance over a distribution class characterized by $\ell_{2}$ separation between sparse cluster centers. Given a variable subset $S$, the clustering step of our algorithms presented in Section 3 solves a reduced version of SDP problem (2) by restricting to variables in $S$:
@@ 2.2 Minimax rate for separation bound in isotropic case
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\mathbf{Z}\rangle\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~{}\mathbf{Z}\succeq 0,~{}\mathrm{tr}(\mathbf{Z})=K,~{}\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~{}\text{and}~{}\mathbf{Z}\geq 0,$ (3)
The analysis proceeds in two steps: First, we show that if there exists a feature subset whose $\ell_{2}$ separation (i.e., $\Delta$ in (4) below) exceeds a threshold (explicitly determined by the sample size $n$, dimension $p$, and sparsity level $s$) then exact cluster label recovery is possible. Second, we construct a worst-case distribution where the separation is no more than a constant multiple of this threshold, under which no method can achieve exact recovery. This establishes that no clustering algorithm can succeed under weaker separation conditions. Together, these results underscore the role of sparsity in problem difficulty and serves as a motivation of our iterative algorithm that incorporates variable selection for addressing sparsity.
@@
We begin by specifying the minimax framework. For each parameter $\boldsymbol{\theta}=(\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{K},\mathbf{H})$, where $\boldsymbol{\mu}_{k}\in\mathbb{R}^{p}$ are cluster centers and $\mathbf{H}\in\{0,1\}^{n\times K}$ denotes the cluster assignment matrix, as explained in (1), let $\mathbb{P}_{\boldsymbol{\theta}}$ be the corresponding joint distribution of $n$ independent Gaussian vectors. Each vector has mean given by the assigned cluster center and covariance $\sigma^{2}\mathbf{I}_{p}$. For all $1\leq k\neq l\leq K$, we impose two restrictions on $\boldsymbol{\mu}_{k}-\boldsymbol{\mu}_{l}$. First, it has at most $s$ nonzero entries. Second, its squared $\ell_{2}$ norm exceeds a specified threshold $\Delta^{2}$. These conditions define the following Gaussian mixture distribution class:

### v2

@@ Abstract
We study high-dimensional clustering where the true signal is sparse among features. To help design algorithms that do not rely on precise estimation of sparse nuisance parameters or signal feature set, we establish minimax separation for exact cluster recovery of semidefinite programming (SDP) relaxation of $K$-means using varying subsets of features. It highlights the critical role of feature selection and, at the same time, shows that the signal feature set can be slightly overestimated without compromising clustering performance. Guided by this theory, our method alternates between rough feature selection and clustering: feature selection is performed by thresholding a rough estimate of the discriminative direction, while clustering is carried out via an SDP-relaxed $K$-means. We further extend the method to settings with unknown sparse precision matrices, avoiding full model parameter estimation by computing only the minimally required quantities.
@@ 1 INTRODUCTION
High-dimensional clustering methods group unlabeled observations when the number of features exceeds or matches the sample size. They are widely applied to high-cost recordings with limited sample sizes such as gene expression (Curtis et al.,, 2012; He et al.,, 2021) and neurophysiological signals (Cunningham and Yu,, 2014; Kadir et al.,, 2014; Park et al.,, 2024). They are also used in machine learning tasks such as reinforcement learning (Liu and Frank,, 2022), privacy protection (Javanmard and Mirrokni,, 2023), and federated learning (Kim et al.,, 2024). In many such problems, only a small subset of features is relevant to the clustering. For example, in document clustering, each document contains only a few informative words out of a large corpus (Bing et al.,, 2020; Wu et al.,, 2023; Yi et al.,, 2014). In disease subtype discovery from gene expression data, only a handful of genes out of thousands differ across subtypes (Golub et al.,, 1999).
If sparsity is present and effectively leveraged, clustering performance improves, as demonstrated in broad applications such as genomics (Lu et al.,, 2017), neuroimaging data analysis (Mishra et al.,, 2017; Zhang et al.,, 2021; Namgung et al.,, 2024), and financial modeling (Nystrup et al.,, 2021). However, sparsity in clustering introduces a key challenge of interdependence between the two unsupervised tasks of feature selection and clustering. On one hand, with unknown sparsity, including too many noise features in the clustering algorithm hinders the results. On the other hand, it is challenging to pinpoint the signal features when the true cluster structure is unknown. This interplay motivates two-step or iterative algorithms that alternates between feature selection and clustering, leading to the identification of the true clusters.
Sparse Clustering Model
We use the following Gaussian mixture model to study the role of sparsity in clustering. Let $\mathbf{X}\in\mathbb{R}^{p\times n}$ be the data matrix with columns $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}$ independently generated as
$\mathbf{X}_{i}\sim\mathcal{N}(\boldsymbol{\mu}^{\ast}_{k},\boldsymbol{\Sigma}^{\ast})~\text{if}~i\in G_{k}^{\ast},$ (1)
where the non-random sets $G_{1}^{\ast},\ldots,G_{K}^{\ast}$ partition $\{1,\ldots,n\}$. The number of clusters $K$ is assumed known. The unknown cluster centers are denoted by $\boldsymbol{\mu}^{\ast}_{1},\ldots,\boldsymbol{\mu}^{\ast}_{K}\in\mathbb{R}^{p}$, and $\boldsymbol{\Sigma}^{\ast}\in\mathbb{R}^{p\times p}$ is the common covariance matrix. We consider both cases where $\boldsymbol{\Sigma}^{\ast}$ is known and where it is unknown but $\boldsymbol{\Omega}^{\ast}:=(\boldsymbol{\Sigma}^{\ast})^{-1}$ has a sparse structure. The goal is to recover $G_{1}^{\ast},\dots,G_{K}^{\ast}$, treating $\boldsymbol{\mu}^{\ast}_{k}$’s and $\boldsymbol{\Sigma}^{\ast}$ as nuisance parameters. We assume the existence of an unknown signal feature set
$S_{0}:=\bigcup_{1\leq k\neq l\leq K}\mathrm{supp}\bigl(\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})\bigr)\subset\{1,\ldots,p\},$ (2)
which encodes the complete clustering signal. This assumption aligns with results from classification of two-class Gaussian mixtures, where the optimal decision boundary maximizing the signal-to-noise ratio is proportional to $\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$(Fan et al.,, 2013).
Previous Two-Step Methods
Conventional dimension reduction, such as PCA (Ghosh and Chinnaiyan,, 2002; Liu et al.,, 2003; Tamayo et al.,, 2007), can fail to capture a sparse clustering signal (Chang,, 1983). Consequently, many methods adopt models such as (1) and (2) to perform feature selection guided by the clustering signal. They select features using multiple testing based on goodness-of-fit (Jin and Wang,, 2016) or moments (Azizyan et al.,, 2013; Verzelen and Arias-Castro,, 2017), model selection (Maugis et al.,, 2009; Raftery and Dean,, 2006), sparse PCA (Löffler et al.,, 2022), merging algorithm (Banerjee et al.,, 2017), or an estimated sparse discriminating direction (defined in Section 3.1; Azizyan et al.,, 2015). Other relevant works include Bouveyron and Brunet-Saumard, (2014); Brodinová et al., (2019); Kadir et al., (2014); Liu et al., (2023). Jin et al., (2017) show that, in certain regimes of signal strength and sparsity, accurate feature selection is not required for accurate clustering. Even et al., (2025) use data thinning (Leiner et al.,, 2025; Dharamshi et al.,, 2025) to decouple feature selection from clustering and achieves optimality among polynomial-time clustering algorithms. They further show that the separation required for the success of their method differs from the minimax separation needed for optimal statistical performance without computational constraints.
Previous Iterative Methods
@@
Iterative methods repeatedly refine cluster assignments and model parameters estimates. Most methods alternate between clustering and regularized estimation of model parameters, such as feature weights (Chakraborty et al.,, 2023; Huo and Tseng,, 2017; Li et al.,, 2018; Tsai and Chiu,, 2008; Witten and Tibshirani,, 2010; Yuan et al.,, 2024), cluster centers (Sun et al.,, 2012; Yi et al.,, 2014), and feature rankings (Arias-Castro and Pu,, 2017; Zhang et al.,, 2020). Regularized EM algorithms follow the same principle: the E-step updates cluster assignments, while the M-step performs regularized estimation of sparse model parameters (Bouveyron et al.,, 2007; Dong et al.,, 2024; Fraley and Raftery,, 2007; Fu et al.,, 2021; Guo et al.,, 2010; Pan and Shen,, 2007; Wang and Zhu,, 2008; Xie et al.,, 2008; Zhou et al.,, 2009). Notably, Cai et al., (2019) establish a minimax rate for the excess clustering risk based on estimating $\boldsymbol{\beta}^{\ast}$ from the training data using $\ell_{1}$-regularized EM. Sparse Bayesian hierarchical models (Tadesse et al.,, 2005; Yao et al.,, 2025) and sparse convex clustering (Chakraborty and and Xu,, 2023; Wang et al.,, 2018) also follow the iterative paradigm, as their fitting procedures, such as MCMC or ADMM, progressively refine parameter estimates and feature selection.
Limitations of Existing Approaches
In each iteration of existing iterative methods, sparse model parameters are first estimated, which then define the clustering optimization problem. Estimation requires tuning sparsity-inducing regularization parameters, which is challenging in unsupervised settings and computationally heavy. However, across the computer science (Wilder et al.,, 2019; Bengio et al.,, 2021), operations research (Bertsimas and Kallus,, 2020; Elmachtoub and Grigas,, 2022; Sadana et al.,, 2025), and statistics (Perdomo et al.,, 2020; Cutler et al.,, 2024) literature, it is recognized that when estimation supports a downstream task, prioritizing task quality over estimation accuracy is often preferable. This motivates an algorithm that avoids precise estimation of nuisance parameters and instead focuses on clustering quality.
Our Contributions
SDP $K$-means is a minimax-optimal clustering algorithm for fixed $p$ that bypasses cluster center estimation (Chen and Yang,, 2021). We quantify its sensitivity to feature selection accuracy by deriving the minimax separation under varying feature subsets. Uniformly over all subsets $S$, SDP $K$-means restricted to $S$ exactly recovers the clusters if the separation with respect to $S\cap S_{0}$, as defined in (8), exceeds $\log n+(|S|\log p)/n+\sqrt{(|S|\log p)/n}.$ We show that it is minimax optimal in the regime where $\tfrac{|S|\log p}{n}\lesssim\log n$. This indicates that slight overestimation of $S_{0}$ is not highly detrimental. Guided by this theory, we propose an iterative algorithm that alternates between feature selection and SDP $K$-means, where feature selection does not require precise estimation of nuisance parameters. Features are roughly selected by thresholding an estimate of the discriminating direction $\boldsymbol{\beta}^{\ast}$ (defined in Section 3.1), which guides $S_{0}$ estimation but does not directly affect clustering. Based on the roughly estimated $S_{0}$, we run SDP $K$-means on a transformed dataset $\boldsymbol{\Omega}^{\ast}\mathbf{X}$ (innovated transformation; Fan and Lv,, 2016), avoiding precise estimation of cluster centers or covariance matrix. To obtain the transformed dataset, we adapt a high-dimensional precision matrix estimation tool (Fan and Lv,, 2016), estimating only the minimally required quantities and bypassing high-dimensional nuisance parameter estimation.
Notations
@@
We denote vectors and matrices by boldface letters, and their individual elements by the corresponding non-boldface letters. For a matrix $\mathbf{A}$, single and double subscripts denote columns and (row, column) elements, respectively. A subscript may be a single index or an index set. For vectors not derived from a matrix, a subscript denotes an element or a sub-vector. If the vector or matrix already contains a subscript, we use parentheses to avoid ambiguity, for example, $(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{S}$. We write $\mathbf{A}\geq 0$, $\mathbf{A}>0$, and $\mathbf{A}\succeq 0$ to denote elementwise nonnegativity, elementwise positivity, and positive semidefiniteness of a matrix, respectively. For a set $S$, let $|S|$ denotes its cardinality. For an integer $a>0$, let $\mathbf{1}_{a}\in\mathbb{R}^{a}$ denote the all-one vector, and $[a]$ denote $\{1,\cdots,a\}$. For a set $G_{k}\subset[n]$, let $\mathbf{1}_{G_{k}}\in\{0,1\}^{n}$ denote the indicator vector that takes the value 1 at indices in $G_{k}$ and 0 elsewhere.
2 THEORETICAL MOTIVATION
We first introduce SDP $K$-Means and highlight its limitations in high dimensions through the lens of minimax analysis.
SDP $K$-means
@@ 2 THEORETICAL MOTIVATION
The partition $G_{1},\ldots,G_{K}$ maximizing the $K$-means objective is the solution of the following NP-hard mixed integer program (Peng and Wei,, 2007; Zhuang et al.,, 2023):
$\max_{\mathbf{H}\in\{0,1\}^{n\times K}}\;\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{H}\mathbf{B}\mathbf{H}^{\top}\rangle,~\text{s.t.}~\mathbf{H}\mathbf{1}_{K}=\mathbf{1}_{n},$ (3)
@@
where the inner product denotes the Frobenius inner product, $\mathbf{B}:=\mathrm{diag}(|G_{1}|,\ldots,|G_{K}|)$, and each row of $\mathbf{H}$ is a one-hot cluster indicator, formally $H_{i,k}:=\mathbbm{1}(\mathbf{X}_{i}\in G_{k})$ for $i\in[n]$ and $k\in[K]$. Appendix E.1 connects the objective (3) to the log likelihood function in the Gaussian mixture model. Since the change of variable $\mathbf{Z}=\mathbf{H}\mathbf{B}\mathbf{H}^{\top}$ satisfies $\mathbf{Z}^{\top}=\mathbf{Z},\mathbf{Z}\succeq 0,\mathrm{tr}(\mathbf{Z})=K,\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n}$, and $\mathbf{Z}\geq 0$, we can lift $\mathbf{Z}$ into the convex space and relax (3) into a semidefinite program:
@@
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}~\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{Z}\rangle\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~\mathbf{Z}\succeq 0,~\mathrm{tr}(\mathbf{Z})=K,~\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~\mathbf{Z}\geq 0.$ (4)
@@
The SDP $K$-means algorithm recovers cluster labels by applying spectral clustering to the solution of (4). When $\boldsymbol{\Sigma}^{\ast}$ is known to be $\mathbf{I}_{p}$ and the dimensionality $p$ is fixed, SDP $K$-means achieves the minimax separation under the exact cluster recovery loss (Chen and Yang,, 2021), where the separation is defined as in (8) with $S=[p]$. The minimax separation $\log n+\sqrt{(\log n)^{2}+(Kp\log n)/n}$ increases with $p$ at a square-root rate. This shows that any clustering method that does not adapt to sparsity, including SDP $K$-means, may fail in high-dimensional settings where $p$ is comparable to or larger than $n$, unless the separation also scales quickly at the square-root rate with $p$ — an assumption that can be very stringent for large $p$. This motivates us to modify the SDP $K$-means to incorporate sparsity in high dimensions and explore the corresponding separation rate.
Minimax Separation under Sparsity
We study how the separation needed for SDP $K$-means to achieve exact recovery scales with sample size, ambient dimension, sparsity and estimation quality of $S_{0}$. We assume that $\boldsymbol{\Sigma}^{\ast}$ is known to be $\sigma^{2}\mathbf{I}_{p}$. Given any variable subset $S$, consider SDP $K$-means restricted to $S$, which solves the SDP problem (4) over the subset $S$:
@@
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}~\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\mathbf{Z}\rangle\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~\mathbf{Z}\succeq 0,~\mathrm{tr}(\mathbf{Z})=K,~\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~\mathbf{Z}\geq 0,$ (5)
and then obtain cluster labels by applying spectral clustering to the solution. The minimax analysis proceeds in two steps. First, we show that if the clustering signal (or separation) in $S\cap S_{0}$, defined in (8), exceeds a certain threshold explicitly depending on the sample size $n$, ambient dimension $p$, and subset size $|S|$, then exact recovery of cluster labels is possible with high probability. This high-probability event is established uniformly over all possible subsets $S$. Second, we identify the minimax optimal regime for our result. Together, these results reveal the role of sparsity in clustering difficulty. They also motivate our iterative algorithm, which incorporates feature selection to address sparsity and solves a sequence of SDPs corresponding to the sequence of selected feature subsets.
@@
We begin by specifying the problem setting. For each parameter tuple $\boldsymbol{\theta}=(\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{K},\mathbf{H})$, with $\boldsymbol{\mu}_{k}\in\mathbb{R}^{p}$ cluster centers and $\mathbf{H}\in\{0,1\}^{n\times K}$ cluster assignment matrix, let $\mathbb{P}_{\boldsymbol{\theta}}$ be the corresponding joint distribution of $n$ independent Gaussian random vectors. Each random vector has mean given by the assigned cluster center and covariance $\sigma^{2}\mathbf{I}_{p}$. For all $1\leq k\neq l\leq K$, we impose two restrictions on $\boldsymbol{\mu}_{k}-\boldsymbol{\mu}_{l}$. First, it has at most $s$ nonzero entries. Second, its squared $\ell_{2}$ norm exceeds a specified threshold $\Delta^{2}$. These conditions define the following Gaussian mixture distribution class: